(x^2 -yz)p + (y^2 -zx)q = z^2 -xy


1
Expert's answer
2021-09-27T16:21:23-0400

"(x^2-yz)p+(y^2-zx)q=z^2-xy\\
p=\frac{\delta z}{\delta x} \text{ and } q=\frac{\delta z}{\delta y}\\
\text{The auxilliary equations are }\\
\frac{dx}{x^2-yz}=\frac{dy}{y^2-zx}=\frac{dz}{z^2-xy}\\
\text{Hence }\\
\frac{dx-dy}{(x^2-yz)-(y^2-zx)}=\frac{dy-dz}{(y^2-zx)-(z^2-xy)}=\frac{dz-dx}{(z^2-xy)-(x^2-yz)}\\
\frac{d(x-y)}{(x-y)(x+y+z)}=\frac{d(y-z)}{(x-y)(x+y+z)}=\frac{d(x-y)}{(z-x)(x+y+z)}\\
\frac{d(x-y)}{x-y}=\frac{d(y-z)}{y-z}=\frac{d(z-x)}{z-x}\\
\text{Integrate both sides}\\
\ln|x-y|=\ln|y-z|+\ln C_1\\
\ln|y-z|=\ln|z-x|+ \ln C_2\\
\implies \frac{x-y}{y-z}=C_1\\
~~~~~~~~~~\frac{y-z}{z-x}=C_2\\
\text{The general solution of the equation is }\\
\phi\left(\frac{x-y}{y-z},\frac{y-z}{z-x}\right)=0"